# VTechAGP: An Academic-to-General-Audience Text Paraphrase Dataset and Benchmark Models

### v1

@@ Abstract
Existing text simplification or paraphrase datasets mainly focus on sentence-level text generation in a general domain. These datasets are typically developed without using domain knowledge. In this paper, we release a novel dataset, VTechAGP, which is the first academic-to-general-audience text paraphrase dataset consisting of 4,938 document-level these and dissertation academic and general-audience abstract pairs from 8 colleges authored over 25 years. We also propose a novel dynamic soft prompt generative language model, DSPT5. For training, we leverage a contrastive-generative loss function to learn the keyword vectors in the dynamic prompt. For inference, we adopt a crowd-sampling decoding strategy at both semantic and structural levels to further select the best output candidate. We evaluate DSPT5 and various state-of-the-art large language models (LLMs) from multiple perspectives. Results demonstrate that the SOTA LLMs does not provide satisfactory outcomes, while the lightweight DSPT5 can achieve competitive results. To the best of our knowledge, we are the first to build a benchmark dataset and solutions for academic-to-general-audience text paraphrase dataset.
@@ 1 Introduction
Existing text simplification and paraphrase generation datasets (shown in Table 5 in the Appendix) mainly focus on sentence-level translation. The recent paragraph-level dataset WikiAuto Jiang et al. (2020), derived from WikiLarge Zhang and Lapata (2017) and Newsela Xu et al. (2015), still suffers from a lack of domain diversity and specification, which are sourced from Wikipedia and news articles. Besides, the objective of these studies on the aforementioned datasets is to simplify the text for children at lower grade levels. Furthermore, existing works on lay summarization involves brief summarization limited to medicine and biological domains Devaraj et al. (2021); Flores et al. (2023); Jiang et al. (2024); Zaman et al. . However, our proposed VTechAGP involves translation from technical language to general-audience language across a broader set of multiple domains. We aim to generate text from the domain-specific to a general level of understanding, while keeping it scientifically accurate and easy to understand, in order to encourage and facilitate interdisciplinary collaboration across different research fields.
In this paper, we pioneer the research of academic-to-general-audience text generation by introducing a new benchmark dataset VTechAGP, which is derived from electronic theses and dissertations (ETDs) at Virginia Tech over twenty-five years. VTechAGP consists of 4,938 document-level abstract pairs (academic abstract and general-audience abstract). VTechAGP also provides other information such as title, discipline, degree level, etc. This auxiliary information shows the potential of VTechAGP for other tasks such as topic generation, etc. In addition, the abstracts in VTechAGP come from multiple domains (colleges) and VTechAGP is labeled with each specific domain and provides the domain knowledge keywords. More details about VTechAGP are presented in Sec. 3 and Table 6.
@@ 2.2 Text Generation Methods
College # Doc. Avg. # Sentence Avg. Sentence Len. MTLD Readability Consensus
Agriculture and Life Sciences (ALS) 429 15.06 / 13.80 27.42 / 26.37 73.59 / 68.58 15th-16th / 14th-15th
Architecture, Arts, and Design (AAD) 226 10.83 / 8.99 26.67 / 26.07 68.24 / 68.76 13th-14th / 10th-11th
Engineering (ENG.) 2451 14.03 / 11.99 26.87 / 25.66 69.93 / 66.47 13th-14th / 14th-15th
Liberal Arts and Human Sciences (LAHS) 675 10.12 / 8.60 30.24 / 28.46 68.02 / 67.06 16th-17th / 15th-16th
Natural Resources and Environment (NRE) 269 14.48 / 12.66 28.87 / 27.33 76.69 / 71.99 15th-16th / 10th-11th
Science (SCI.) 679 14.19 / 11.75 27.35 / 25.73 74.90 / 68.48 16th-17th / 15th-16th
Business (BUS.) 63 11.76 / 10.24 28.55 / 27.63 68.46 / 70.35 18th-19th / 17th-18th
Veterinary Medicine (VM) 146 15.61 / 13.16 26.83 / 25.92 77.77 / 70.15 16th-17th / 14th-15th
All 4938 13.68 / 11.66 27.68 / 26.38 71.36 / 67.91 15th-16th / 13th-14th
@@ 3.1 Data Source and Dataset Collection
Virginia Tech has been a leader in electronic theses and dissertations (ETDs) for more than twenty-five years. It was the first university to require electronic submission of ETDs, beginning in 1997. Virginia Tech’s ETDs are accessible through VTechWorks, [^1] https://vtechworks.lib.vt.edu/ a digital repository created through a collaboration between the Graduate School and the University Libraries. In the fall of 2016, the Graduate School added a new requirement for ETDs: the inclusion of a general audience abstract in addition to the traditional academic abstract. The ETD submission system was updated in 2019 to include a separate field for the general audience abstract. Since then, most ETDs include both an academic abstract and a general audience abstract, which are captured as distinct metadata fields in VTechWorks. Like many institutional repositories, VTechWorks supports the Open Archives Initiative Protocol for Metadata Harvesting (OAI-PMH) Lagoze et al. (2002). OAI-PMH is a standard protocol for retrieving metadata records from digital repositories. It provides a framework that enables data exchange between various systems, supporting consistent metadata formats. To create the VTechAGP dataset, we used OAI-PMH to harvest metadata from VTechWorks. We specifically queried the VTechWorks OAI-PMH endpoint to retrieve metadata records for ETDs containing both academic and general audience abstracts. The OAI-PMH endpoint provided us with an XML record for each ETD. This data encompassed a range of metadata elements defined in a qualified Dublin Core schema. We identified specific fields necessary for our dataset and extracted the text content. This step involved parsing the XML structure, locating the relevant elements, and retrieving the textual data they contained. We mapped the extracted metadata to specific columns in a CSV file. Each piece of extracted metadata was mapped to a corresponding column header in the CSV file: ‘identifier_uri’, ‘title’, ‘abstract’, ‘abstract_general’, ‘subject_terms’, ‘discipline’, ‘department’, ‘degree’, ‘degree_level’, and ‘type’. A description is given in Table 6 in the Appendix. We provide the data analysis in Sec. B.
@@ 3.2 Task Definition
Figure 1: Dynamic soft prompt generation of our proposed model DSPT5. The framework includes a prompt encoder and a generative language model fine-tuning together with a hybrid loss.
We name the dataset VTechAGP and propose the academic to the general audience text paraphrase task. Let $D=\left\{X_{n},Y_{n},A_{n}\right\}_{n=1}^{N}$ denote a dataset where $X_{n}$ is the source document, $Y_{n}$ is the target document, $A_{n}$ is the auxiliary information (subject terms, which are the keywords in theses/dissertations provided by the author) of the source document, and $N$ is the number of documents. Given an input sequence of words $X=[x_{1},\cdots,x_{S}]$ with length $S$ and auxiliary information $A=[a_{1},\cdots,a_{A}]$ with length $A$, we aim to generate an output sequence of words $Y=[y_{1},\cdots,y_{T}]$ with length $T$ that retains the original meaning as $X$, but reduces the complexity to improve readability and comprehension without domain knowledge. In this task, our goal is to find the set of model parameters that maximize $\prod_{n=1}^{N}p_{model}(Y_{n}|X_{n})$.
@@ 4.2.1 Dynamic Prompt Generation
To better control the output generated by T5 for our task, we added a prompt template before the source input (academic abstracts). Instead of using a static prompt template, such as "Generate another version of the provided document for general audiences", we design a dynamic prompt template that can learn to generate academic keywords to adapt to different academic domains. As opposed to adding the keywords directly in front of the prompt Blinova et al. (2023), we designed a template (which is used in the experiments): "Generate the document by reducing the domain knowledge of " + keywords, where the keywords (auxiliary information) $K^{pos}=[k^{pos}_{1},\cdots,k^{pos}_{N}]$ are generated by KeyBERT Grootendorst (2020) 333 https://github.com/MaartenGr/KeyBERT and sorted by their importance score:
@@ 4.2.3 Model Training with Hybrid Loss
where $y_{n,s}$ is the s-th word of the general audience abstract in the n-th sample, $S_{n}$ is the length of the target output $y_{n}$, and $\theta$ gives the parameters of DSPT5. To further improve the performance of DSPT5, we add an additional contrastive loss that forces the prompt encoder to generate the embeddings of extracted keywords from KeyBERT $r^{pos}$ that can be more similar to the representations of subject terms $r^{key}$ provided from the source data, while steering away from negative extracted words $r^{neg}$, to encourage the model to bring $r^{key}$ and $r^{neg}$ closer in the learned feature space, while pushing dissimilar instances $r^{key}$ and $r^{neg}$ apart. Therefore, we propose a new hybrid loss function consisting of the cross-entropy loss in Equ. 6 and a contrastive loss infoNCE van den Oord et al. (2018) $\mathcal{L}_{nce}$:
$\mathcal{L}_{nce}=-log\frac{exp(\frac{r^{key}\cdot r^{pos}}{\tau})}{exp(\frac{r^{key}\cdot r^{pos}}{\tau})+\sum_{r^{neg}}^{N^{neg}}exp(\frac{r^{key}\cdot r^{neg}}{\tau})}$ (7)
@@ 5.2 Baselines and Evaluation Metrics
We compare DSPT5 with the following LLM baselines: Pre-trained LLaMA2 Touvron et al. (2023)  [^4] We implement the 7B version of LLaMA2., ChatGPT Brown et al. (2020)  [^5] https://chatgpt.com/, and Claude2 Anthropic (2023); Fine-tuned BART Lewis et al. (2020), T5 Raffel et al. (2020), and FLAN-T5 Chung et al. (2022). Crowd sampling is only added to DSPT5 for evaluation, as it is one component of DSPT5. All other LLMs baselines use their default decoding strategies. We evaluate DSPT5 with other SOTA LLMs using the following automatic evaluation metrics. (1) Embedding-based metrics: BERTScore (F1) Zhang* et al. (2020), document-level BLONDE (F1) Jiang et al. (2022), sentence-level and document-level BLEU Lin and Och (2004); (2) Word-based metrics: ROUGE1, ROUGE2 Lin (2004) and METEOR Banerjee and Lavie (2005); (3) End-to-end metrics: COMET Rei et al. (2022); (4) Simplicity: SARI Xu et al. (2016a).(5) Diversity: LCTR (Lexical Translation Consistency Ratio) Lyu et al. (2021).(6) Readability: FRES (Flesch Reading-Ease Score) Flesch (1979); (7) Toxicity Team et al. (2022) for safety evaluation.
@@ 5.3.1 Main Results
Table 2: Performance evaluation of all baselines with twelve different metrics. The abbreviations stands for Agriculture and Life Science, Architecture, Arts and Design, Engineering, Liberal Arts and Human Sciences, Natural Resources and Environment, Science, Business, and Veterinary Medicine.
@@
Table 3: Ablation study over DSPT5 components on VTechAGP across eight colleges. DSPT5dynamic is the model only using dynamic hard prompt from keyBERT. DSPT5dyn+soft is the model using dynamic soft prompt from the prompt encoder. DSPT5dyn+soft+con is the model fine-tuned using dynamic soft prompt with contrastive learning loss. DSPT5all is the final version of the model using dynamic soft prompt with contrastive learning for training and crowd sampling for inference.
@@
DSPT5dynamic 8.78 22.83 83.83 28.43 49.77 25.33 38.35 77.28 36.71 60.19 32.31 0.24
DSPT5dyn+soft 9.87 25.79 84.43 34.92 51.14 26.98 40.64 78.91 37.03 59.11 32.34 0.11
DSPT5dyn+soft+con 9.88 26.78 84.58 33.99 51.66 26.82 40.90 78.54 37.03 59.05 32.33 0.13
DSPT5all 9.93 26.99 84.52 35.75 51.45 27.59 41.40 79.06 37.05 58.77 32.39 0.17
@@ 5.3.3 Human Evaluation
Table 4: Mean human evaluation ratings 1-5 (the higher the better) of different models on VTechAGP. Reported from left to right are: comprehensiveness, layness, meaning preservation, conciseness and fluency.
COM LAY MP CON FLU
BART 3.83 2.83 3.67 3.58 3.83
T5 3.61 2.95 3.35 3.5 3.51
FlanT5 3.71 3.13 3.41 3.66 3.76
Claude2 3.08 2.95 2.72 3.55 3.26
ChatGPT 4.17 3.66 3.83 3.71 4.31
LLaMA2 4.05 3.61 3.73 3.80 4.31
Ours 4.17 3.57 3.90 3.87 3.95
The human evaluation setup is discussed in detail in Sec. F in the Appendix. Human evaluation results are shown in Table 4. From Table 4, we observe that generally ChatGPT, and our proposed DSPT5 show better performance than other baselines. Based on the observations, new information with simpler words is always introduced in the generated outputs from ChatGPT. Those sentences with simpler words make the generated abstracts easier to understand, resulting in a better layness. However, it also significantly changes the meaning of the original sentences and makes the sentences longer with redundency. That’s why ChatGPT shows the best performance in layness, while not the best in meaning preservation and conciseness. DSPT5 attempts to paraphrase the abstracts with non-technical words while retaining the original meaning as closely as possible. While keeping the content concise, it may sacrifice fluency between sentences. Although ChatGPT shows competitive results, it is NOT an open-sourced model, which may raise concerns about data security and high cost issues. Alternatively, the DSPT5 can be fully controlled and owned with only a single V100 GPU for fine-tuning jobs for the academic-to-general-audience text paraphrasing task.

### v2

@@ Abstract
Existing text simplification or paraphrase datasets mainly focus on sentence-level text generation in a general domain. These datasets are typically developed without using domain knowledge. In this paper, we release a novel dataset, VTechAGP, which is the first academic-to-general-audience text paraphrase dataset consisting of document-level thesis and dissertation academic and general-audience abstract pairs from 8 colleges authored over 25 years. We also propose a novel dynamic soft prompt generative language model, DSPT5, for the academic-to-general-audience text paraphrasing task. For training, we leverage a contrastive-generative loss function to learn the keyword vectors in the dynamic prompt. For inference, we adopt a crowd-sampling decoding strategy at both semantic and structural levels to further select the best output candidate. We evaluate DSPT5 and various state-of-the-art large language models (LLMs) from multiple perspectives. Results demonstrate that the SOTA LLMs do not provide satisfactory outcomes, while the lightweight DSPT5 can achieve competitive results. To the best of our knowledge, we are the first to build a benchmark dataset and solutions for academic-to-general-audience text paraphrase dataset  [^1] https://github.com/waingram/VTechAGP-Dataset, Dataset: https://doi.org/10.5281/zenodo.14833932.
@@ 1 Introduction
Existing text simplification and paraphrase generation datasets (shown in Table 5 in the Appendix) mainly focus on sentence-level translation. The recent paragraph-level dataset WikiAuto Jiang et al. (2020), derived from WikiLarge Zhang and Lapata (2017) and Newsela Xu et al. (2015), still suffers from a lack of domain diversity and specification, which are sourced from Wikipedia and news articles. Besides, the objective of these studies on the aforementioned datasets is to simplify the text for children at lower grade levels. Furthermore, existing works on lay summarization involve brief summarization limited to medicine and biological domains Devaraj et al. (2021); Flores et al. (2023); Jiang et al. (2024); Zaman et al. . However, our proposed VTechAGP involves translation from technical language to general-audience language across a broader set of multiple domains. We aim to generate text from the domain-specific to a general level of understanding while keeping it scientifically accurate and easy to understand, in order to encourage and facilitate interdisciplinary collaboration across different research fields.
In this paper, we pioneer the research of academic-to-general-audience text generation by introducing a new benchmark dataset VTechAGP, which is derived from electronic theses and dissertations (ETDs) at Virginia Tech over twenty-five years. VTechAGP consists of document-level abstract pairs (academic abstract and general-audience abstract). VTechAGP also provides other information such as title, discipline, degree level, etc. This auxiliary information shows the potential of VTechAGP for other tasks such as topic generation, etc. In addition, the abstracts in VTechAGP come from multiple domains (colleges) and VTechAGP is labeled with each specific domain and provides the domain knowledge keywords. More details about VTechAGP are presented in Sec. 3 and Table 6.
@@ 2.2 Text Generation Methods
College Avg. # Sentence Avg. Sentence Len. MTLD Readability Consensus
Agriculture and Life Sciences (ALS) 15.06 / 13.80 27.42 / 26.37 73.59 / 68.58 15th-16th / 14th-15th
Architecture, Arts, and Design (AAD) 10.83 / 8.99 26.67 / 26.07 68.24 / 68.76 13th-14th / 10th-11th
Engineering (ENG.) 14.03 / 11.99 26.87 / 25.66 69.93 / 66.47 13th-14th / 14th-15th
Liberal Arts and Human Sciences (LAHS) 10.12 / 8.60 30.24 / 28.46 68.02 / 67.06 16th-17th / 15th-16th
Natural Resources and Environment (NRE) 14.48 / 12.66 28.87 / 27.33 76.69 / 71.99 15th-16th / 10th-11th
Science (SCI.) 14.19 / 11.75 27.35 / 25.73 74.90 / 68.48 16th-17th / 15th-16th
Business (BUS.) 11.76 / 10.24 28.55 / 27.63 68.46 / 70.35 18th-19th / 17th-18th
Veterinary Medicine (VM) 15.61 / 13.16 26.83 / 25.92 77.77 / 70.15 16th-17th / 14th-15th
All 13.68 / 11.66 27.68 / 26.38 71.36 / 67.91 15th-16th / 13th-14th
@@ 3.1 Data Source and Dataset Collection
Virginia Tech has been a leader in ETDs for more than twenty-five years. It was the first university to require electronic submission of ETDs, beginning in 1997. Virginia Tech’s ETDs are accessible through VTechWorks, [^2] Virginia Tech Electronic Theses and Dissertations: https://hdl.handle.net/10919/5534 a digital repository created through a collaboration between the Graduate School and the University Libraries. In the fall of 2016, the Graduate School added a new requirement for ETDs: the inclusion of a general audience abstract in addition to the traditional academic abstract. The ETD submission system was updated in 2019 to include a separate field for the general audience abstract. Since then, most ETDs have included both an academic abstract and a general audience abstract, which are captured as distinct metadata fields in VTechWorks. Like many institutional repositories, VTechWorks supports the Open Archives Initiative Protocol for Metadata Harvesting (OAI-PMH) Lagoze et al. (2002). OAI-PMH is a standard protocol for retrieving metadata records from digital repositories. It provides a framework that enables data exchange between various systems, supporting consistent metadata formats. To create the VTechAGP dataset, we used OAI-PMH to harvest metadata from VTechWorks. We specifically queried the VTechWorks OAI-PMH endpoint to retrieve metadata records for ETDs containing both academic and general audience abstracts. The OAI-PMH endpoint provided us with an XML record for each ETD. This data encompassed a range of metadata elements defined in a qualified Dublin Core schema. We identified specific fields necessary for our dataset and extracted the text content. This step involved parsing the XML structure, locating the relevant elements, and retrieving the textual data they contained. We mapped the extracted metadata to specific columns in a CSV file. Each piece of extracted metadata was mapped to a corresponding column header in the CSV file: ‘identifier_uri’, ‘title’, ‘abstract’, ‘abstract_general’, ‘subject_terms’, ‘discipline’, ‘department’, ‘degree’, ‘degree_level’, and ‘type’. A description is given in Table 6 in the Appendix. We provide the data analysis in Sec. B.
@@ 3.2 Task Definition
Figure 1: Dynamic soft prompt generation of our proposed model DSPT5. The framework includes a prompt encoder and a generative language model fine-tuning together with a hybrid loss. FC is a fully connected layer.
We name the dataset VTechAGP and propose the academic to the general audience text paraphrase task. Let $D=\left\{X_{n},Y_{n},A_{n}\right\}_{n=1}^{N}$ denote a dataset where $X_{n}$ is the source document, $Y_{n}$ is the target document, $A_{n}$ is the auxiliary information (subject terms, which are the keywords in ETDs provided by the author) of the source document, and $N$ is the number of documents. Given an input sequence of words $X=[x_{1},\cdots,x_{S}]$ with length $S$ and auxiliary information $A=[a_{1},\cdots,a_{A}]$ with length $A$, we aim to generate an output sequence of words $Y=[y_{1},\cdots,y_{T}]$ with length $T$ that retains the original meaning as $X$, but reduces the complexity to improve readability and comprehension without domain knowledge. In this task, our goal is to find the set of model parameters that maximize $\prod_{n=1}^{N}p_{model}(Y_{n}|X_{n})$.
@@ 4.2.1 Dynamic Prompt Generation
To better control the output generated by T5 for our task, we added a prompt template before the source input (academic abstracts). Instead of using a static prompt template, such as "Generate another version of the provided document for general audiences", we design a dynamic prompt template that can learn to generate academic keywords to adapt to different academic domains. As opposed to adding the keywords directly in front of the prompt Blinova et al. (2023), we designed a template (which is used in the experiments): "Generate the document by reducing the domain knowledge of " + keywords, where the keywords (auxiliary information) $K^{pos}=[k^{pos}_{1},\cdots,k^{pos}_{N}]$ are generated by KeyBERT Grootendorst (2020) 444 https://github.com/MaartenGr/KeyBERT and sorted by their importance score:
@@ 4.2.3 Model Training with Hybrid Loss
where $y_{n,s}$ is the s-th word of the general audience abstract in the n-th sample, $S_{n}$ is the length of the target output $y_{n}$, and $\theta$ gives the parameters of DSPT5. To further improve the performance of DSPT5, we add an additional contrastive loss that forces the prompt encoder to generate the embeddings of extracted keywords from KeyBERT $r^{pos}$ that can be more similar to the representations of subject terms $r^{key}$ provided from the source data, while steering away from negative extracted words $r^{neg}$, to encourage the model to bring $r^{key}$ and $r^{neg}$ closer in the learned feature space while pushing dissimilar instances $r^{key}$ and $r^{neg}$ apart. Therefore, we propose a new hybrid loss function consisting of the cross-entropy loss in Equ. 6 and a contrastive loss infoNCE van den Oord et al. (2018) $\mathcal{L}_{nce}$:
$\mathcal{L}_{nce}=-log\frac{exp(\frac{r^{key}\cdot r^{pos}}{\tau})}{exp(\frac{r^{key}\cdot r^{pos}}{\tau})+\sum_{r^{neg}\in N^{neg}}exp(\frac{r^{key}\cdot r^{neg}}{\tau})}$ (7)
@@ 5.2 Baselines and Evaluation Metrics
We compare DSPT5 with the following LLM baselines: Pre-trained LLaMA2 Touvron et al. (2023)  [^5] We implement the 7B version of LLaMA2., ChatGPT Brown et al. (2020)  [^6] https://chatgpt.com/, and Claude2 Anthropic (2023); Fine-tuned BART Lewis et al. (2020), T5 Raffel et al. (2020), and FLAN-T5 Chung et al. (2022). Each baseline is fine-tuned on each domain. Crowd sampling is only added to DSPT5 for evaluation, as it is one component of DSPT5. All other LLM baselines use their default decoding strategies. We evaluate DSPT5 with other SOTA LLMs using the following automatic evaluation metrics. (1) Embedding-based metrics: BERTScore (F1) Zhang* et al. (2020), document-level BLONDE (F1) Jiang et al. (2022), sentence-level and document-level BLEU Lin and Och (2004); (2) Word-based metrics: ROUGE1, ROUGE2 Lin (2004) and METEOR Banerjee and Lavie (2005); (3) End-to-end metrics: COMET Rei et al. (2022); (4) Simplicity: SARI Xu et al. (2016a).(5) Diversity: LCTR (Lexical Translation Consistency Ratio) Lyu et al. (2021).(6) Readability: FRES (Flesch Reading-Ease Score) Flesch (1979); (7) Toxicity Team et al. (2022) for safety evaluation.
@@ 5.3.1 Main Results
Table 2: Performance evaluation of all baselines with twelve different metrics. The abbreviations stand for Agriculture and Life Science, Architecture, Arts and Design, Engineering, Liberal Arts and Human Sciences, Natural Resources and Environment, Science, Business, and Veterinary Medicine.
@@
Table 3: Ablation study over DSPT5 components on VTechAGP across eight colleges. DSPT5dynamic is the model only using a dynamic hard prompt from keyBERT. DSPT5dyn+soft is the model using a dynamic soft prompt from the prompt encoder. DSPT5dyn+soft+con is the model fine-tuned using dynamic soft prompt with contrastive learning loss. DSPT5all is the final version of the model using dynamic soft prompt with contrastive learning for training and crowd sampling for inference.
@@
Oursdynamic 8.78 22.83 83.83 28.43 49.77 25.33 38.35 77.28 36.71 60.19 32.31 0.24
Oursdyn+soft 9.87 25.79 84.43 34.92 51.14 26.98 40.64 78.91 37.03 59.11 32.34 0.11
Oursdyn+soft+con 9.88 26.78 84.58 33.99 51.66 26.82 40.90 78.54 37.03 59.05 32.33 0.13
Oursall 9.93 26.99 84.52 35.75 51.45 27.59 41.40 79.06 37.05 58.77 32.39 0.17
@@ 5.3.3 Human Evaluation
Table 4: Mean human evaluation ratings 1-5 (the higher the better) of different models on VTechAGP. Reported from left to right are: comprehensiveness, layness, meaning preservation, conciseness, and fluency. ICC stands for the intraclass correlation coefficient.
COM LAY MP CON FLU ICC
BART 3.83 2.83 3.67 3.58 3.83 0.85
T5 3.61 2.95 3.35 3.5 3.51 0.60
FlanT5 3.71 3.13 3.41 3.66 3.76 0.38
Claude2 3.08 2.95 2.72 3.55 3.26 0.67
ChatGPT 4.17 3.66 3.83 3.71 4.31 0.85
LLaMA2 4.05 3.61 3.73 3.80 4.31 0.39
Ours 4.17 3.57 3.90 3.87 3.95 0.78
Following a similar setting as Liu et al. (2024); Li et al. (2024); Song et al. (2024); Kew et al. (2023); Devaraj et al. (2021), our evaluation uses a random sample of 20 abstracts from the test split VTechAGP considering the workload. Judges are presented with both the academic abstract and generated general-audience abstracts from seven models for each data sample in a total of 140 abstracts. Using a 1-5 Likert scale, the judges are asked to rate the model output based on five criteria: comprehensiveness, layness, meaning preservation, conciseness, and fluency. More details are discussed in detail in Sec. F in the Appendix.
Human evaluation results are shown in Table 4. From Table 4, we observe that generally ChatGPT, and our proposed DSPT5 show better performance than other baselines. ChatGPT performs well in comprehensiveness, layness, and fluency, whereas DSPT5 outperforms all other baselines in comprehensiveness, meaning perservation, and conciseness. Based on the observations, new information with simpler words is always introduced in the generated outputs from ChatGPT. Those sentences with simpler words make the generated abstracts easier to understand, resulting in better layness. However, it also significantly changes the meaning of the original sentences and makes the sentences longer with redundancy. That’s why ChatGPT shows the best performance in layness, while not the best in meaning preservation and conciseness. DSPT5 attempts to paraphrase the abstracts with non-technical words while retaining the original meaning as closely as possible. While keeping the content concise, it may sacrifice fluency between sentences.
For intraclass correlation coefficient (ICC)  [^7] We use Pingouin package to calculate ICC: https://pingouin-stats.org/build/html/index.html, which is used to determine if items or subjects can be rated reliably by different raters, BART, ChatGPT and DSPT5 show good reliability. T5 and Claude2 show moderate reliability. FlanT5 and LLaMA2 show poor reliability according to Koo and Li (2016). Although ChatGPT shows competitive results, it is NOT an open-sourced model, which may raise concerns about data security and high-cost issues. Alternatively, the DSPT5 can be fully controlled and owned with only a single V100 GPU for fine-tuning jobs for the academic-to-general-audience text paraphrasing task.